OpenStudy (anonymous):

solve the equation 1 over 4 x + 2 = 5 for x. 28 20 14 12

OpenStudy (anonymous):

i am going to make a bet that this is \[\frac{1}{4}x+2=5\] am i right?

OpenStudy (anonymous):

yess just gotta solve it

OpenStudy (anonymous):

I assumed that as well. So what you do is you first isolate the variable. So 1/4 x = 3. See how I did it? Then you would divide 3 by 1/4. And get x

OpenStudy (anonymous):

is it 12?

OpenStudy (anonymous):

what @SavannahWillett said subtract 2 from both sides to get \[\frac{1}{4}x=3\] but i would say "multiply both sides by 4" to get \[x=3\times 4\] rather than "divide by \(\frac{1}{4}\)" but it is really the same thing

OpenStudy (anonymous):

yes, it is 12